A wave is given by the equation y(x, t) = 4.0cos(0.5πx−10 πt) m. The speed of the wave is:
1
Expert's answer
2020-04-17T10:16:10-0400

The equation of a wave

"y(x,t)=A \cos\left(\frac{2\pi}{T} t-\frac{2\pi}{\lambda}x\right)"

Hence, "T=4\:\rm s, \; \lambda=0.2\:\rm m."

The speed of wave

"v=\frac{\lambda}{T}=\frac{0.2}{4}=0.05\:\rm m/s=5\:\rm cm/s."